### v1

# Hidden Markov Model Filtering with Equal Exit Probabilities

Abstract
Hidden Markov Models (HMMs) provide a rigorous framework for inference in dynamic environments. In this work, we study the $\alpha$-HMM algorithm motivated by the optimal online filtering formulation in settings where the true state evolves as a Markov chain with equal exit probabilities. We quantify the dynamics of the algorithm in stationary environments, revealing a trade-off between inference and adaptation, showing how key parameters and the quality of observations affect performance. Comprehensive theoretical analysis on the nonlinear dynamical system that governs the evolution of the log-belief ratio over time and numerical experiments demonstrate that the proposed approach effectively balances adaptation and inference performance.
I Introduction
In this work, we consider the online filtering problem, where an agent receives noisy observations or signals $\boldsymbol{\xi}_{i}$ arising from some time-varying state $\boldsymbol{\theta}_{i}^{\star}$ at each time step $i=1,2,\ldots$. The aim is to estimate the underlying true state $\boldsymbol{\theta}_{i}^{\star}$ at each time instant $i$ given the observations $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{i}$.
For tractability, we assume that the true state $\boldsymbol{\theta}_{i}^{\star}$ belongs to a discrete set of $M$ possible states $\Theta=\{\theta_{0},\theta_{1},\ldots,\theta_{M-1}\}$. The agent assigns a belief to each state $\theta\in\Theta$ at each time step $i$, denoted by $\boldsymbol{\mu}_{i}(\theta)$, which characterizes the agent’s confidence that $\theta$ is the underlying true state at time $i$ based on previous observations, i.e.
$\boldsymbol{\mu}_{i}(\theta)=\mathbb{P}\left[\boldsymbol{\theta}_{i}^{\star}=\theta|\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{i}\right],\quad\theta\in\Theta.$ (1)
Correct learning is said to occur at time $i$ if the belief $\boldsymbol{\mu}_{i}(\theta)$ is maximized at the true state $\boldsymbol{\theta}_{i}^{\star}$. Conditioned on $\boldsymbol{\theta}_{i}^{\star}$, the observations are independent random variables over time $i$, taking values in the space $\Xi$. Given the underlying true state $\theta_{i}^{\star}$, the observations follow a probability density function $f(\cdot|\theta_{i}^{\star})$, which implies that when the underlying state of the environment remains unchanged, the observations are independent and identically distributed (i.i.d.) random variables over time. The agent is equipped with a model that specifies the likelihood of the observations $\xi$ for each possible state $\theta\in\Theta$, denoted by
$L(\xi|\theta),\quad\xi\in\Xi.$ (2)
In a static environment, where the underlying true state remains constant over time, the exact belief (1) can be computed exactly and recursively using Bayes’ formula, resulting in the following update rule for each possible state indexed by $m$:
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\boldsymbol{\mu}_{i-1}(\theta_{m})L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}\boldsymbol{\mu}_{i-1}(\theta_{n})L(\boldsymbol{\xi}_{i}|\theta_{n})}.$ (3)
Such constructions are widely applied in decision-making, learning, and inference tasks for single or multi-agent systems in static environments. In multi-agent networks or social networks, the sequential process where agents infer the true state by observing the decisions of peers and signals generated by the environment, is referred to as (non-)Bayesian social learning[1, 2, 3, 4, 5, 6, 7, 8]. These models are widely used in economics, political science, and sociology to model the behavior of financial markets, social groups, and social networks[9, 10].
In dynamic environments, where the true state evolves over time, the reliance of (3) on past observations can make it less effective. In such scenarios, optimal filtering methods include the Kalman filter[11] and the hidden Markov filter[12], with the latter assuming that the evolution of the true state follows a Markov chain. In this paper, we focus on filtering models related to the hidden Markov framework. Assuming that the transition probability matrix of the Markov chain between the $M$ states at each iteration is denoted as $P=\left[p_{nm}\right]_{M\times M}$, where $p_{nm}=\mathbb{P}\left[\theta_{i}^{\star}=\theta_{m}\mid\theta_{i-1}^{\star}=\theta_{n}\right]$, then the optimal belief update, based on the hidden Markov model (HMM), is given by:
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\sum_{n=0}^{M-1}p_{nm}\boldsymbol{\mu}_{i-1}(\theta_{n})L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{l=0}^{M-1}\sum_{n=0}^{M-1}p_{nl}\boldsymbol{\mu}_{i-1}(\theta_{n})L(\boldsymbol{\xi}_{i}|\theta_{l})}.$ (4)
Observe that the optimal belief update under a hidden Markov model involves the entries of the full state transition matrix $P$. In practice, $P$ is frequently unknown and challenging to estimate. In this work, we will instead study a simplified HMM-based update, which is derived under the assumption of equal exit probabilities for state transitions. Specifically, the true state transitions with a probability $h$, and when a transition occurs, the next state is chosen uniformly at random. Under these assumptions, the transition probability matrix simplifies to:
$p_{mn}=\begin{cases}1-h,&m=n,\\ \frac{h}{M-1},&m\neq n.\end{cases}$ (5)
This transforms the traditional HMM belief update into:
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{((1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha)L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}((1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{n})+\alpha)L(\boldsymbol{\xi}_{i}|\theta_{n})},$ (6)
where $\alpha=\frac{h}{M-1}$ represents the exit probability. This formulation, referred to as the $\alpha$-HMM, simplifies the filtering problem to a single tunable hyperparameter $\alpha$. Of course, for a general state transition matrix $P$, the simplified update rule (6) will be suboptimal compared to the optimal HMM filter (4). The advantage, on the other hand, is that (6) relies only on a single paramater $\alpha$, that quantifies the volatility of the underlying true state. Indeed, examining (6), we observe that $\alpha$ essentially controls the amount of weight placed on prior beliefs compared to the most recent observation. Note that when $\alpha=0$, the above iteration degenerates into the classical Bayes’ update (3). In this paper, since we focus on filtering in dynamic environments, we only consider the case where $10<h<$ 1 and consequently, $\alpha>0$.
In the sequel, we analyze the steady-state behavior of the $\alpha$-HMM algorithm (6) and quantify the tradeoff between learning and adaptability introduced by the hyperparameter $\alpha$. Numerical experiments validate our theoretical findings and highlight the practical applicability of the $\alpha$-HMM in dynamic environments.
I-A Related Work
The hidden Markov model, proposed in the late 1960s[13], remains widely used due to its ease of implementation, ability to handle sequential data, and capacity to process inputs of variable lengths, making it suitable for numerous real-life applications[14]. Applications of HMMs are extensively studied in fields such as speech recognition[15, 16, 17], human activity recognition[18, 19, 20], and data analysis[21].
Theoretical studies on the dynamical analysis of HMMs often focus on the stability of the algorithm, particularly the conditions under which a wrongly initialized belief converges to the true posterior distribution. Filter stability, as discussed in [22], arises through two primary mechanisms: the ergodicity of the transition model[23, 24] and the informativeness of the observations[25].
The ergodicity of the transition model, which has been more extensively studied, is often analyzed through the Dobrushin coefficient. The Dobrushin coefficient maps the transition probabilities of the Markov chain to a real number in the interval $\left[0,1\right]$ and determines how quickly the HMM filter forgets its initial conditions and converges to the stationary distribution. In [24], a novel sufficient condition for the exponential stability of HMMs is proposed, leveraging Dobrushin coefficients associated with both the transition kernel and the informativeness of the observations.
In this work, we adopt a dynamical systems perspective, utilizing concepts such as fixed points to analyze the stochastic and deterministic dynamics of an HMM with equal exit probabilities. The ergodicity of the transition model can be easily adjusted by tuning the exit probability parameter $\alpha$, and the informativeness of the observations is incorporated through an identifiability parameter.
The online filtering problem has also been extensively studied in distributed settings. The adaptive social learning (ASL) strategy introduced in [26] is a linear strategy based on the logarithmic likelihood ratio of beliefs. ASL provides strong analytical guarantees and exhibits effective learning and adaptability in dynamic environments. Building on this framework, [8] implements multi-agent collaborative online machine learning. A distributed HMM framework for multi-agent systems was proposed in [27], enabling agents to collaboratively track dynamic states governed by Markov chains. Furthermore, [28] evaluates the performance of ASL by modeling the true state as a rare transition Markov chain.
I-B Notations
Throughout this paper, we use $i$ to denote time iteration, bold notation to represent random quantities, while $\mathbb{P}\left[\cdot\right]$ and $\mathbb{E}\left[\cdot\right]$ represent the probability and expectation operators, respectively.
We denote the $L_{p}$-norm of a vector by $\left\|\cdot\right\|_{p}$, i.e.,
$\left\|v\right\|_{p}=\left(\sum_{i=1}^{N}\left|v_{i}\right|^{p}\right)^{1/p},$ (7)
where $v=\left[v_{1},\ldots,v_{N}\right]^{\top}$, $0<p<\infty$, and
$\left\|v\right\|_{\infty}=\underset{i=1,\ldots,N}{\max}\left|v_{i}\right|.$ (8)
We use $0_{N}$ to denote the $N$-dimensional column vectors with all elements equal to 0.
II Steady State Dynamical Analysis
Assume the environment remains in a single state over an extended period, resulting in fixed observation statistics. Without loss of generality, we assume that the underlying true state is $\theta_{0}\in\Theta$, i.e. $\theta_{i}^{\star}=\theta_{0}$ for all $i=1,2,\ldots$. We aim to analyze the dynamics of the log-belief ratio, i.e.:
$\boldsymbol{x}_{m,i}\triangleq\log\frac{\boldsymbol{\mu}_{i}(\theta_{m})}{\boldsymbol{\mu}_{i}(\theta_{0})},\quad m=1,\ldots,M-1,\quad i=0,1,\ldots.$ (9)
It can be verified that following (6), the log-belief ratio evolves as:
$\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})+\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})},$ (10)
where
$F_{m}(x_{1},\ldots,x_{M-1})\triangleq$ (11)
$\log\frac{(1-\alpha M)\exp(x_{m})+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}.$
To ensure the agent can successfully learn the underlying true state, we impose the following assumptions:
Assumption 1 (Finiteness of KL Divergence)
For each pair of distinct states $\theta$ and $\theta^{\prime}$, the Kullback–Leibler (KL) divergence[29] between $L(\xi|\theta)$ and $L(\xi|\theta^{\prime})$ satisfies:
$D_{KL}(L(\xi|\theta)||L(\xi|\theta^{\prime}))<\infty.$ (12)
This assumption avoids trivial cases where a likelihood function model for a certain state completely dominates.
Assumption 2 (Identifiability of the Underlying True State)
$\left\{\theta_{i}^{\star}\right\}=\Theta_{i}^{\star}\triangleq\underset{\theta\in\Theta}{\arg\min}D_{\text{KL}}(f(\cdot|\theta_{i}^{\star})||L(\cdot|\theta)).$ (13)
Assumption 2 ensures that the underlying true state is the single best match for the observations under the agent’s likelihood model, which guarantees that the agent can achieve successful learning using its likelihood model and the observed data.
Remark 1
The nonlinear function (11) has the following properties:
$F_{m}(0,\ldots,0)=0,$ (14)
$\frac{\partial F_{m}}{\partial x_{m}}\Big{|}_{(0,\ldots,0)}=1-\alpha M,$ (15)
$\frac{\partial F_{m}}{\partial x_{n}}\Big{|}_{(0,\ldots,0)}=0,\quad\forall n\neq m.$ (16)
By applying a multivariate Taylor expansion to $F(x)$ around $x=0_{M-1}$ up to the first-order term, for all $m=1,\ldots,M-1$ we have:
$F_{m}(x_{1},\ldots,x_{M-1})=(1-\alpha M)x_{m}+o(\left\|x\right\|),$ (17)
then we derive the linear approximation of the system (10), which is:
$\boldsymbol{x}_{m,i}=(1-\alpha M)\boldsymbol{x}_{m,i-1}+\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (18)
The above equation can be seen as a model that resembles the single-agent version of the Adaptive Social Learning model [26] in non-Bayesian social learning, which has the following dynamics for log-belief ratio:
$\boldsymbol{x}_{m,i}=(1-\delta)\boldsymbol{x}_{m,i-1}+\delta\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})},$ (19)
where $\delta$ is the step-size parameter. Equations (18) and (19) both apply a discount to the information from the previous time step during the iteration process. The subtle difference between them lies in the fact that the latter normalizes the weighting between past and new information through the step-size parameter $\delta$.
It can be seen that the recursion (10) is both nonlinear and stochasticnonlinear since the first term of the r.h.s. is a nonlinear function of $\boldsymbol{x}_{m,i-1}$ and stochastic since the system is driven by the function of a random process $\boldsymbol{\xi}_{i}$.
To facilitate analysis, we introduce the following deterministic reference dynamical system, where stochastic quantities are replaced by their expected values:
$\hat{x}_{m,i}=F_{m}(\hat{x}_{1,i-1},\ldots,\hat{x}_{M-1,i-1})-d_{m},$ (20)
where $m=1,\ldots.M-1$ and
$d_{m}$ $\triangleq D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{m}))-D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{0}))$ (21)
$=-\mathbb{E}\left[\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}\right].$
Assumption 2 guarantees that $d_{m}>0$. From the definition of $d_{m}$, it is evident that it characterizes the agent’s ability to distinguish any incorrect state $\theta_{m}$ from the underlying true state $\theta_{0}$ using the likelihood model and the observation $\boldsymbol{\xi}_{i}$. A larger $d_{m}$ indicates stronger identifiability.
In this paper, we first analyze the $(M-1)$-dimensional deterministic dynamical system given by (20) as a reference system. Subsequently, we will demonstrate that due to the contraction properties of the function $F(\cdot)$, the asymptotic difference between the original stochastic system (10) and this reference system can be bounded.
In the following theoretical analysis, we focus on the scenario where $0<\alpha<1/M$, ensuring $10<1-\alpha M<$ 1. This condition, referred to as slow-changing conditions, not only aligns better with intuitive expectations—since a high value of $\alpha$ (indicative of highly volatile environments) is expected to degrade the algorithm’s performance—but also facilitates a more analytically tractable analysis of the algorithm.
II-A Convergence Rate and Adaption Ability
In this part, we will prove that the reference dynamical system (20) will converge to a fixed point at an exponential rate. Based on this, we can analyze the adaptation capability and how it can be influenced by the exit probability parameter $\alpha$ and $d_{m}$.
The following theorem provides the main result regarding the convergence of the reference dynamical system (20) and its convergence rate.
Theorem 1
When $0<\alpha<1/M$, $\hat{x}_{i}$ defined in (20) converges to a unique fixed point, i.e.,
$\lim_{i\rightarrow\infty}\hat{x}_{i}=\hat{x}^{\infty},$ (22)
where
$\hat{x}_{i}=\left[\hat{x}_{1,i},\ldots,\hat{x}_{M-1,i}\right]^{\top},$ (23)
the fixed point is given by
$\hat{x}^{\infty}=\left[\hat{x}_{1}^{\infty},\ldots,\hat{x}_{M-1}^{\infty}\right]^{\top},$ (24)
$\hat{x}^{\infty}_{m}=\log\frac{\alpha}{\alpha\exp(d_{m})+(1-\alpha M)(\exp(d_{m})-1)\hat{\mu}_{0}^{\infty}},$ (25)
$\hat{\mu}_{0}^{\infty}=\frac{1}{1+\sum_{m=1}^{M-1}\exp(\hat{x}_{m}^{\infty})}.$ (26)
Furthermore, the convergence rate is bounded by:
$\left\|\hat{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\leq\lambda\left\|\hat{x}_{i-1}-\hat{x}^{\infty}\right\|_{\infty},$ (27)
where the contraction factor is given by:
$\lambda=1-\min\left\{\alpha,\frac{\underline{d}}{\overline{x}_{0}-\log\frac{\alpha}{1-\alpha M+\alpha}+\underline{d}}\right\},$ (28)
$\overline{x}_{0}=\max\left\{\left\{\hat{x}_{m,0}\right\}_{m=1}^{M-1},0\right\},$ (29)
and
$\underline{d}=\underset{m=1,\ldots,M-1}{\min}d_{m}.$ (30)
Proof:
Proof omitted due to space limitations. ∎
Remark 2
Theorem 1 shows that the reference dynamical system converges to a fixed point, which is strictly less than zero for all components. To see this, for all $m=1,\ldots,M-1$, when $01-\alpha M>$ 0 we have:
$\alpha\exp(d_{m})+(1-\alpha M)(\exp(d_{m})-1)\hat{\mu}^{\infty}_{0}>\alpha\exp(d_{m}),$ (31)
implying
$x_{m}^{\infty}<-d_{m}<0.$ (32)
This result means that in the reference dynamical system, the agent places higher belief on the underlying true state at the fixed point. When $01-\alpha M>$ 0, the higher the value of $d_{m}$, which represents the agent’s ability to distinguish between $\theta_{0}$ and $\theta_{m}$, the more confident the agent is in the underlying true state at the fixed point.
Remark 3
Theorem 1 provides an upper bound on the convergence rate of $\hat{x}_{i}$ to $\hat{x}^{\infty}$ in the sense of the $L_{\infty}$-norm, which decreases as the exit probability $\alpha$ increases, as shown in (28). We can also observe the impact of the data quality, captured in $\underline{d}$, and the initial setting, captured in $\overline{x}_{0}$, on the convergence rate of the reference dynamical system to its fixed point. The parameter $\underline{d}$ depends on both the local likelihood function (e.g., pre-trained models) and the quality of observations (e.g., variance of local data noise). This parameter, referred to as the identifiability parameter, reflects the agent’s ability to distinguish between the two most challenging states. From Theorem 1, it follows that higher data quality—manifested as a larger $\underline{d}$—accelerates the convergence to the fixed point, while the deviation in the initial belief, captured in $\overline{x}_{0}$, slows down the convergence rate.
The impact of the exit probability $\alpha$ and the identifiability parameter $\underline{d}$ on the convergence rate can be further illustrated by the following lemma.
Corollary 1
When $0<\alpha<1/M$, the convergence rate of the reference dynamical system has the following property:
$\left\|\exp(\hat{x}_{i})-\exp(\hat{x}^{\infty})\right\|_{1}\leq\gamma\left\|\exp(\hat{x}_{i-1})-\exp(\hat{x}^{\infty})\right\|_{1},$ (33)
where
$\gamma=\frac{(1-\alpha M)\left(1+2\alpha U(M-1)\right)}{(1-\alpha M+\alpha)^{2}}\exp(-\underline{d}),$ (34)
$U=\max\left\{1,\left\{\exp(\hat{x}_{m,0})\right\}_{m=1}^{M-1}\right\},$ (35)
$\underline{d}=\underset{m=1,\ldots,M-1}{\min}d_{m}.$ (36)
Furthermore, if the initial value of the reference dynamical system satisfies:
$\hat{x}_{0}\geq\hat{x}^{\infty},$ (37)
then a preciser bound of contraction rate can be given by:
$\left\|\exp(\hat{x}_{i})-\exp(\hat{x}^{\infty})\right\|_{1}\leq\gamma^{\prime}\left\|\exp(\hat{x}_{i-1})-\exp(\hat{x}^{\infty})\right\|_{1},$ (38)
where
$\gamma^{\prime}=\frac{\exp(-\underline{d})(1-\alpha M)}{(1-\alpha M+\alpha)^{2}}.$ (39)
Proof:
Proof omitted due to space limitations. ∎
Corollary 1 further shows that as $\alpha\rightarrow 1/M$, contraction coefficient $\gamma\rightarrow 0$, indicating that the rate of convergence accelerates. This observation aligns with intuition: when $\alpha\rightarrow 1/M$ in equation (6), the system relies less on historical observations and becomes more reliant on the most recently observed data. Additionally, as $\underline{d}\rightarrow\infty$, $\gamma\rightarrow 0$, which again shows that higher data quality leads to a faster convergence rate of the system to its fixed point.
When the components of the initial values of the reference dynamical system satisfy certain consistency conditions, a more precise upper bound on the convergence rate can be provided. It is worth noting that when the algorithm adopts a uniform prior belief distribution:
$\boldsymbol{\mu}_{0}(\theta_{m})=\boldsymbol{\mu}_{0}(\theta_{n}),\quad\forall\theta_{m},\theta_{n}\in\Theta,$ (40)
we have $\hat{x}_{0}=0_{M-1}$. The initial condition (37), will therefore be satisfied. This condition reflects an unbiased prior belief, which aligns with typical requirements in practical algorithmic settings.
Remark 4
The convergence rate of the reference dynamical system to its fixed point, as presented in Theorem 1, plays a crucial role in understanding the adaptation capabilities of the algorithm. Specifically, when the environment remains in a fixed state for an extended period, the reference dynamical system converges to its fixed point at an exponential rate. When the environment state changes, the corresponding fixed point shifts accordingly, and the system will converge to the new fixed point exponentially fast, as illustrated in Fig. 1. The convergence speed depends on the choice of the parameter $\alpha$ and the quality of the data and model. When using the traditional Bayes’ formula (3), as indicated by equations (11) and (20) with $\alpha=0$, the corresponding deterministic dynamical system for the log-belief ratio reduces to a linear equation, which evolves linearly over time. This slower rate explains why the proposed algorithm exhibits superior adaptability in dynamic environments.
Figure 1: (a) The agent decides between two possible states, $\theta_{0}$(state 1) and $\theta_{1}$(state 2), where the likelihood models for these two states, as well as the model for environmental observations under these two states, are both standard normal density functions with distinct means. The environment remains in state 1 until time $T_{1}$, after which it switches to state 2 for the remaining duration. (b) The log-belief ratio $\log(\boldsymbol{\mu}_{i}(\theta_{1})/\boldsymbol{\mu}_{i}(\theta_{0}))$ is compared over time for both the $\alpha$-HMM and the traditional Bayes’ filter. Here $T^{A}_{\text{adapt}}$ and $T^{B}_{\text{adapt}}$ represent the adaptation time—specifically, the time required for the log-belief ratio to shift from negative to positive—required for the $\alpha$-HMM and Bayes’ formula, respectively. The HMM filter quickly converges and stabilizes at the fixed point, whereas the Bayes’ filter continues to accumulate ”confidence” in the current state. When the environment changes, the Bayes’ filter must counteract the previously accumulated confidence and adjust at a slower, linear rate. This slower adjustment prevents the Bayes’ filter from rapidly responding to changes in the true state. The difference in adaptability between the two algorithms is particularly pronounced when $T_{1}$ is large. (c) The belief on the true state under both algorithms over time. In deterministic expected dynamics, the $\alpha$-HMM exhibits superior adaptability, with the belief remaining above 0.5 for most of the time, indicating correct decision-making. When observation noise is present, the $\alpha-$HMM algorithm exhibits larger fluctuations to maintain adaptability. This trade-off allows the $\alpha$-HMM to remain responsive, albeit at the cost of increased variance in belief estimates.
II-B Fixed Point Value and Learning Performance
In this section, we analyze the value of the fixed point. Since we have already proven that the reference dynamical system converges to the unique fixed point as described in Theorem 1, the position of the fixed point will be sufficient to characterize the resulting learning performance. Due to the nonlinearity of (25), explicitly solving for the exact fixed point value is challenging. Nevertheless, we can provide an estimation of its value. In Theorem 2, we derive the lower bound for the counterpart of the agent’s belief $\boldsymbol{\mu}_{i}(\cdot)$ in the reference dynamical system.
Given that
$\boldsymbol{x}_{m,i}=\log\frac{\boldsymbol{\mu}_{i}(\theta_{m})}{\boldsymbol{\mu}_{i}(\theta_{0})},$ (41)
we can obtain
$\boldsymbol{\mu}_{i}(\theta_{m})=\begin{cases}1/\left(1+\sum_{m=1}^{M-1}\exp(\boldsymbol{x}_{m,i})\right),&m=0,\\ \exp(\boldsymbol{x}_{m,i})/\left(1+\sum_{m=1}^{M-1}\exp(\boldsymbol{x}_{m,i})\right),&m\neq 0.\end{cases}$ (42)
Since $\hat{x}_{m,i}$ can be seen as the counterpart of $\boldsymbol{x}_{m,i}$ in reference dynamical system, similarly we can define
$\hat{\mu}_{i}(\theta_{m})=\begin{cases}1/\left(1+\sum_{m=1}^{M-1}\exp(\hat{x}_{m,i})\right),&m=0,\\ \exp(\hat{x}_{m,i})/\left(1+\sum_{m=1}^{M-1}\exp(\hat{x}_{m,i})\right),&m\neq 0,\end{cases}$ (43)
and then
$\hat{x}_{m,i}=\log\frac{\hat{\mu}_{i}(\theta_{m})}{\hat{\mu}_{i}(\theta_{0})},$ (44)
for all $m=1,\ldots,M-1$ and $i=0,1,\ldots$. Here $\hat{\mu}_{i}(\theta)$ can be seen as the counterpart of agent’s belief on state $\theta$ in the reference dynamical system, and we have the following result:
$\lim_{i\to\infty}\hat{\mu}_{i}(\theta_{m})=\hat{\mu}_{m}^{\infty},\quad\forall m=0,\ldots,M-1,$ (45)
where
$\hat{\mu}_{m}^{\infty}=\begin{cases}1/\left(1+\sum_{m=1}^{M-1}\exp(\hat{x}_{m}^{\infty})\right),&m=0,\\ \exp(\hat{x}_{m}^{\infty})/\left(1+\sum_{m=1}^{M-1}\exp(\hat{x}_{m}^{\infty})\right),&m\neq 0.\end{cases}$ (46)
This result follows directly from Theorem 1 and the properties of limits.
The value of $\hat{\mu}_{0}^{\infty}$, which represents the steady-state belief on the true state in the reference dynamical system, reflects the learning performance of the algorithm: the closer $\hat{\mu}_{0}^{\infty}$ is to 1, the better the algorithm distinguishes between the correct and incorrect hypotheses at steady state; conversely, the closer $\hat{\mu}_{0}^{\infty}$ is to $1/M$, the more ambiguous the agent’s learning outcome. We provide upper and lower bound estimates for $\hat{\mu}_{0}^{\infty}$ in the following lemma.
Lemma 1
When $0<\alpha<1/M$ and the underlying true state remains unchanged, the $\hat{\mu}_{0}^{\infty}$ defined in (46) satisfies the following lower bound:
$\hat{\mu}_{0}^{\infty}\geq\frac{(1-\alpha M)\exp(\underline{d})-1+\alpha}{(1-\alpha M)(\exp(\underline{d})-1)},$ (47)
where
$\underline{d}=\underset{m=1,\ldots,M-1}{\min}d_{m}.$ (48)
Proof:
Proof omitted due to space limitations. ∎
Remark 5
We can observe from Lemma 2 that as $\alpha\to 0$ or $\underline{d}\to\infty$, the steady-state belief on the correct state in the reference dynamical system approaches one, indicating improved learning performance. This observation aligns with intuition: better learning performance is expected in less volatile environments, with lower observation noise, and a more accurate local likelihood model. Combining the result in Lemma 2 with the fixed point values in (25) and (26), we can conclude that as $\alpha\to 0$ or $\underline{d}\to\infty$, $\hat{\mu}_{0}^{\infty}\to 1$ and the fixed point value $\hat{x}_{m}^{\infty}\to-\infty$, which means that the agent becomes more confident in trusting the underlying true state at steady state.
III From Deterministic to Stochastic Systems
In this section, we discuss the relationship between the stochastic dynamical system induced by the $\alpha$-HMM, represented by (10), and our reference deterministic dynamical system (20). In Lemma 2, we show that the expected distance between $\boldsymbol{x}_{i}$ and the fixed point of the reference dynamical system can be asymptotically bounded. Then, in Theorem 2, we compute the instantaneous error probability of the $\alpha$-HMM based on the fixed point of the reference dynamical system. To ensure the feasibility of theoretical analysis, we introduce the following assumption.
Assumption 3 (Bounded Log-Likelihood Ratio)
There exists a positive constant $C$ such that:
$\underset{m=1,\ldots,M-1}{\max}\>\underset{\boldsymbol{\xi}\in\Xi}{\sup}\>\left|\log\frac{L(\boldsymbol{\xi}|\theta_{m})}{L(\boldsymbol{\xi}|\theta_{0})}+d_{m}\right|\leq C,\quad a.s..$ (49)
Note that
$\mathbb{E}\left[\log\frac{L(\boldsymbol{\xi}|\theta_{m})}{L(\boldsymbol{\xi}|\theta_{0})}\right]=-d_{m}.$ (50)
The Assumption 3 actually requires that the log-likelihood ratio can be almost surely bounded.
To compare the reference dynamical system and the original stochastic dynamical system, in the following we analyze two recursions:
1. The evolution of the log-belief ratio in $\alpha$-HMM, $\boldsymbol{x}_{m,i}$ ($m=1,\ldots,M-1$):
$\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})+\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (51)
2. The evolution of the counterpart of $\boldsymbol{x}_{m,i}$ in reference dynamical system, denoted by $\hat{x}_{m,i}$:
$\hat{x}_{m,i}=F_{m}(\hat{x}_{1,i-1},\ldots,\hat{x}_{M-1,i-1})-d_{m},$ (52)
where it is established in Theorem 1 that $\hat{x}_{m,i}\to\hat{x}^{\infty}_{m}$, as $i\to\infty$.
The following lemma establishes the relationship between the two dynamical systems described above:
Lemma 2
Let $\boldsymbol{x}_{m,i}$ and $\hat{x}_{m,i}$ follow the dynamical system in (51) and (52), respectively. Then, the following equation holds:
$\mathbb{E}\left[\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\right]\leq O(\lambda_{1}^{i})+\frac{C}{1-\lambda_{1}},$ (53)
where $\hat{x}^{\infty}$ is as defined in (24),(25) and (26),
$\lambda_{1}=1-\min\left\{\alpha,\frac{\underline{d}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+C}\right\},$ (54)
with $\underline{d}$ and $C$ defined in (30) and (49) respectively.
Proof:
Proof omitted due to space limitations. ∎
In Lemma 2, we show that the distance between the original dynamical system and the fixed point of the reference dynamical system can be bounded. The coefficient $\lambda_{1}$ is always less than 1, then as the time $i\to\infty$, the upper bound of the expected distance between the two is related to the first absolute moment of the log-likelihood ratio of the observations and the contraction coefficient $\lambda_{1}$. In other words, when the observation noise decreases, the stochastic dynamical system corresponding to the $\alpha$-HMM (10) approaches the fixed point of the reference dynamical system.
To evaluate the learning performance of $\alpha$-HMM, we introduce the instantaneous error probability:
$p^{e}_{i}\triangleq\mathbb{P}\left[\;\exists\;\theta_{m}\neq\theta_{0},\text{ s.t. }\boldsymbol{\mu}_{i}(\theta_{m})\geq\boldsymbol{\mu}_{i}(\theta_{0})\right].$ (55)
From this definition, it is clear that the instantaneous error probability quantifies the probability of the agent failing to correctly identify the underlying true state at a given time $i$ during the online filtering process. We will then provide an estimate of this metric in the steady-state scenario for the proposed $\alpha$-HMM algorithm.
Theorem 2
When $0<\alpha<1/M$ and the underlying true state remains constant, i.e. $\theta_{i}^{\star}=\theta_{0}$ for all $i=1,2,\ldots$, the instantaneous error probability for the $\alpha-$HMM algorithm (6) satisfies:
$p_{i}^{e}\leq\frac{O(\lambda_{1}^{i})+C}{-(1-\lambda_{1})\overline{x}^{\infty}},$ (56)
where
$\overline{x}^{\infty}=\underset{m=1,\ldots,M-1}{\max}\hat{x}_{m}^{\infty}<0,$ (57)
other variable definitions remain unchanged from those presented in Lemma 2.
Proof:
Proof omitted due to space limitations. ∎
Remark 6
The contraction coefficient $\lambda_{1}$ in Theorem 3 is always less than 1, consequently the upper bound of the instantaneous error probability decays exponentially with time and eventually converges to a fixed value
$\frac{C}{-(1-\lambda_{1})\overline{x}^{\infty}}.$ (58)
This result indicates that under the $\alpha$-HMM framework, the probability of erroneous learning cannot asymptotically approach zero, even in steady-state conditions, which contrasts with the behavior predicted by Bayes’ formula [7]. However, by sacrificing some learning performance, the $\alpha$-HMM framework significantly enhances adaptability, as previously discussed. The steady-state error learning probability can be reduced either by (1) pushing the equilibrium point $\hat{x}^{\infty}$ of the deterministic system further away from zero and accelerating the convergence rate (i.e., reducing $\lambda_{1}$), noting that, as shown in earlier analyses, improving the identifiability parameter $\underline{d}$ facilitates this outcome, or by (2) reducing the observation noise, thereby increasing the informativeness of the observations.
IV Numerical Experiments
IV-A Performance of $\alpha-$HMM
In the first numerical experiment, we consider a filtering task where the agent observes noisy signals following a standard normal distribution. The number of possible states is set to $M=5$, and the observations follow $\boldsymbol{\xi}_{i}\sim\mathcal{N}(\theta_{i}^{\star},\sigma^{2})$. The agent’s local likelihood function model is given by $L(\boldsymbol{\xi}|\theta_{m})=\mathcal{N}(\theta_{m},\sigma^{2})$, where $\theta_{m}=m+1$, $m=0,\ldots,M-1$. We assume that every ten iterations, the underlying true state, $\theta_{i}^{\star}$, will be randomly reselected from the $M$ possible states. Under this setting, the evolution of the true state does not strictly follow a Markov chain. However, for the filtering problem in such a dynamic environment, our algorithm can still be employed to track the underlying true state.
Figure 2: (a) The likelihood model of the agent. (b) The underlying true state in each iteration. (c-e) Evolution of the belief on the true state for both $\alpha$-HMM and linearized $\alpha$-HMM under different values of $\alpha$ with $\sigma=0.5$ fixed. The success or failure of identification at different iterations is shown, i.e., whether the true state is assigned the highest belief. The black line represents the results for $\alpha$-HMM, the blue line represents the results for linearized $\alpha$-HMM, and the green numbers indicate the success rate of learning over 100 iterations. (f) Top: Learning rates for correctly identifying the true state of the two algorithms under various values of $\alpha$, evaluated via Monte Carlo simulation over 10,000 iterations. Bottom: Evolution of the belief on the true state for both algorithms under their optimal $\alpha$ settings.
Fig. 2 compares the performance of the $\alpha$-HMM algorithm proposed in this paper with the algorithm employing linear log-belief ratio dynamics (18):
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\boldsymbol{\mu}^{1-\alpha M}_{i-1}(\theta_{m})L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}\boldsymbol{\mu}^{1-\alpha M}_{i-1}(\theta_{n})L(\boldsymbol{\xi}_{i}|\theta_{n})},$ (59)
referred to as the linearized $\alpha$-HMM algorithm. It is noteworthy that the belief update rule in (59) exhibits similarities to the one presented in [26], as elaborated in Remark 1. From Fig. 2, it can be observed that under various parameter settings, the accuracy of the $\alpha$-HMM algorithm is no worse than that of the linearized algorithm, and its performance with respect to $\alpha$ is more stable. A more pronounced advantage is observed when $\alpha$ is small. These results highlight the benefits of the nonlinearity inherent in the evolution of the log-belief ratio in $\alpha$-HMM algorithm.
Focusing solely on the $\alpha$-HMM algorithm, when $\alpha$ is relatively small (Fig. 2(c)), the algorithm struggles to track transitions between two similar underlying true states (e.g., at iteration 40, when the true state transitions from $\theta_{2}$ to $\theta_{3}$, the algorithm fails to detect this change). However, as $\alpha$ increases, as shown in Fig. 2(d), the belief on the current true state decreases, which aligns with the analysis results in Lemma 1, allowing the algorithm to better track transitions between similar states, though with some delay. When $\alpha$ is further increased, as illustrated in Fig. 2(e), the algorithm quickly adapts to state transitions, demonstrating excellent adaptability. Nevertheless, due to the reduced confidence in the current true state, the algorithm becomes more sensitive to noise, leading to intermittent correct learning even during stable periods of the true state.
Figure 3: The evolution of the instantaneous error probability with iterations under different parameter settings.
IV-B Simulation of Error Probability
In this example, we simulate the error probability of the $\alpha$-HMM algorithm via numerical experiments. We consider a discrete sampling distribution, where the likelihood models of the agents are assumed to belong to a family of binomial distributions with $n=5$ trials and a probability of success given by $p_{\theta}=\sigma\theta$, where $\theta\in\Theta=\{1,2,3\}$. These distributions are characterized by the following probability mass function:
$L(\xi|\theta)=\binom{n}{\xi}p_{\theta}^{\xi}(1-p_{\theta})^{n-\xi},\quad\xi\in\left\{0,1,\ldots,n\right\},$ (60)
where $\binom{n}{\xi}=\frac{n!}{\xi!(n-\xi)!}$. This model setting can be verified to satisfy Assumption 3.
Fig. 3 illustrates the time evolution of the instantaneous error probability defined in (55) under different parameter settings, as estimated through 10,000 Monte Carlo simulations. A clear trend of decay over time is observed, which is consistent with the predictions of Theorem 2. Additionally, when $\sigma$ is fixed and, consequently, $d_{m}$ remains unchanged, increasing $\alpha$ reduces the contraction coefficient $\lambda_{1}$, thereby accelerating the convergence of the error probability. However, it also shifts the fixed point closer to zero, increasing the probability of error occurrence, as described in (56). Conversely, when $\alpha$ is fixed, an increase in $\sigma$ enhances the agent’s ability to distinguish between states, leading to a larger $d_{m}$. This, in turn, results in a smaller contraction coefficient $\lambda_{1}$, further accelerating the convergence of the error probability. Simultaneously, the fixed point moves further away from zero, thereby reducing the probability of error occurrence. These observations further validate the theoretical analysis presented in Theorem 2.
V Conclusion
In this paper, we studied the online optimal filtering problem by assuming that the environmental state follows a Markov chain with equal exit probability and motivated the $\alpha$-HMM algorithm. By conducting a dynamical analysis of the discrete-time stochastic system associated with the algorithm and its deterministic counterpart under steady-state conditions, we investigated the influence of the exit probability $\alpha$ and the agent’s identifiability on the algorithm’s adaptability and learning performance.
Our findings reveal that a larger $\alpha$ enhances convergence speed but lowers confidence in the correct state at the fixed point, increasing the likelihood of erroneous learning. Thus, the choice of $\alpha$ embodies a tradeoff between adaptability and learning precision. Conversely, higher identifiability—representing superior observation quality or a more accurate local model—accelerates convergence, improves adaptability, and shifts the fixed point further from zero, enhancing overall learning performance.